OpenStudy (anonymous):

4x + y + 14 x-2y=8 Can you explain how to do the problem please?

OpenStudy (mathstudent55):

Is the second + sign in the first equation correct, or should it be an = sign?

OpenStudy (anonymous):

oh it should be an equal sign my bad

OpenStudy (mathstudent55):

Ok. You have a system of equations that you need to solve.

OpenStudy (mathstudent55):

This is the correct system of equations, right? 4x + y = 14 x - 2y = 8

OpenStudy (anonymous):

is it 4,-2?

OpenStudy (mathstudent55):

There are several methods of solving systems of equations. Have you been asked to use a specific method, or is it up to you?

OpenStudy (anonymous):

You can solve the problem by substitution. 4x + y=14 x-2y=8 so first isolate y or x x-2y=8 add 2y to both sides x=2y+8 then substitue 2y+8 into 4x+y=14 4(2y+8)+y=14 distribute 8y+32+y=14 combine like terms 9y+32=14 subtract 32 9y=-18 divide by 9 y=-2 then u plug in -2 in the y into the other equation x=2y+8 x=2(-2)+8 x=-4+8 x=5 so the answer is x=5 y=-2 (5,-2)

OpenStudy (mathstudent55):

Yes, the solution you asked about is correct.

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

@EndlessTrip Great explanation, but are you sure the second line below is correct? x=-4+8 x=5

OpenStudy (mathstudent55):

No. You did everything correctly, and almost at the end, you added -4 + 8 incorrectly. -4 + 8 = 4, not 5. The answer is (4, -2). Look at your original solution, five lines up form the end. That's all. It's a minor arithmetic error.

OpenStudy (anonymous):

Ohhhh lol i didn't even notice. So yeah the answer is (4,-2).

OpenStudy (mathstudent55):

Great. That's much better.

OpenStudy (mathstudent55):

You did a great job explaining the substitution method.